# MixLM: High-Throughput and Effective LLM Ranking via Text-Embedding Mix-Interaction

### v1

@@ Abstract.
Large language models (LLMs) excel at capturing semantic nuances and therefore show impressive relevance ranking performance in modern recommendation and search systems. However, they suffer from high computational overhead under industrial latency and throughput requirements. In particular, cross-encoder ranking systems often create long context prefill-heavy workloads, as the model has to be presented with the user, query and item information. To this end, we propose MixLM, a novel LLM-based ranking framework, which significantly improves the system throughput via reducing the input context length, while preserving the semantic strength of cross-encoder rankers. In contrast to a standard ranking system where the context is presented to the model as pure text, we propose to use mix-interaction, a mixture of text and embedding tokens to represent the input. Specifically, MixLM encodes all items in the catalog into a few embedding tokens and stores in a nearline cache. The encoded item descriptions are used during online inference, effectively reducing the item length from a few thousand text tokens to a few embedding tokens. We share insights from deploying our MixLM framework to a real-world search application at LinkedIn, including a detailed discussion of our training pipelines, as well as a thorough analysis of our online serving infrastructure optimization. Comparing with strong baselines, MixLM increased throughput by 10.0x under the same latency budget, while maintaining relevance metrics. The efficiency gains delivered by MixLM enabled full-traffic deployment of LLM-powered search, which resulted in a significant 0.47% increase in Daily Active Users (DAU) in online A/B tests.
copyright: acmlicensedjournalyear: 2025conference: Preprint; 2025; Arxiv11footnotetext: Equal key contributors.22footnotetext: Work done while at LinkedIn.33footnotetext: Correspondence authors.
@@ 1. Introduction
Large Language Models (LLMs) have been widely adopted across industrial applications due to their strong capabilities in both generative and predictive tasks. In particular, search and recommender systems have emerged as major application areas where LLMs deliver substantial quality improvements over existing approaches (relevance2; relevance3; relevance4; llmrs). For example, when used as a relevance judge, an LLM can interpret subtle linguistic cues and capture conceptual relationships, enabling it to retrieve and rank semantically relevant results that traditional lexical methods may overlook.
Despite their impressive text-understanding capabilities, deploying large language models (LLMs) in search and recommender systems remains challenging due to their substantial computational overhead, particularly when processing long text inputs, making it difficult to meet the stringent industrial requirements for high throughput and low latency. As an example, a standard (cross-encoder) approach to using LLMs for ranking and search applications is to form an input text prompt, including user’s query and candidate item(s) description, that instructs the LLM to determine if the query and the item are relevant (relevance2; relevance3; relevance-biencoder; rankvicuna). This requires the LLM to process long input prompts. Such workloads with heavy prefill often lead to large latency, making them unacceptable for real-world industrial systems, as a result of the quadratic complexity of the self-attention mechanism, a key component of LLM architectures, with regards to the input context length. Often, in real-world implementation of LLM-based semantic search and recommender systems, the LLM has to be replaced with a smaller model, or the features that lead to long input prompts have to be dropped from the online serving stack (related-ebay; related-pinterest; related-walmart). Thus, reducing the input context length for LLM-based ranking and search applications remains an important research problem.
@@
In this work, we introduce MixLM, an LLM-based ranking framework designed to substantially reduce input context length for semantic search applications while preserving the semantic richness of full-text. MixLM operates by compressing each candidate document, often containing thousands of tokens, into a compact set of learned embedding tokens using an LLM encoder. These embeddings are stored in a nearline cache and, during online inference, are mixed with the user’s natural language query (and auxiliary textual signals) and processed by a ranker LLM. By reducing the document representation from thousands of tokens to a handful of embedding tokens, this mixed-input design achieves orders-of-magnitude lower computational complexity, while retaining the rich query–document interactions characteristic of full-text LLM rankers.
@@
Beyond the core modeling approach, we present insights from developing and deploying MixLM in LinkedIn’s production semantic job search system. We describe how to jointly train the encoder and ranker models in an end-to-end manner, and we introduce new distillation losses that effectively transfer knowledge from an existing full-text LLM ranker into the mixed-interaction architecture, ensuring alignment between embedding-based and text-based ranking behaviors. We further study scaling properties of MixLM with respect to training data size and the number of embedding tokens per item.
@@
From a systems perspective, we provide a comprehensive analysis of MixLM’s impact on serving infrastructure. By reducing the context length through embedding compression, MixLM yields substantially higher ranking throughput. We detail optimization techniques tailored for mixed-input LLMs, including nearline caching and amortized prefill computation, and show how these techniques compound the efficiency gains.
@@
Finally, we demonstrate MixLM’s effectiveness at industrial scale through its deployment in LinkedIn’s AI-driven Job Search. MixLM:
@@
Combined with its efficiency advantages, MixLM enabled full LLM-ranking-based search deployment and delivered a significant 0.47% lift in Daily Active Users (DAU) in a large-scale online A/B experiment.
Summary of Contributions:
  • We present MixLM, an industrial-scale ranking framework that aims to reduce the prompt length for ranker LLMs by substituting candidate items’ textual description with a few tokens of embeddings.
  • We introduce an end-to-end training pipeline that jointly optimizes the encoder LLM and the ranker LLM, that allow for capturing rich query–item–feature interactions beyond the capacity of bi-encoder architectures.
  • We describe a production-ready serving stack that incorporates several system optimizations pertaining to our mixed input ranking system, achieving an order-of-magnitude improvement in latency and GPU throughput.
  • We share lessons and insights from developing and deploying our MixLM framework online to LinkedIn’s semantic job search, the primary job search experience on the platform.
Taken together, these advances make LLM-based ranking practically deployable and economically sustainable on the industrial scale, enabling search and recommender systems that demand both deep language understanding and real-time responsiveness.
@@ 2. Semantic Search
Semantic Search
refers to retrieving and ranking items based on their semantic similarity to a given query rather than exact lexical overlap. A typical pipeline includes (1) representation of queries and items into dense embeddings using neural encoders, (2) retrieval of top-$k$ candidates via exhaustive vector search—for example, GPU-accelerated brute-force search, (3) reranking of the retrieved candidates using a higher-capacity model that jointly considers query–item interactions, and (4) application of business logic such as personalization and diversity.
In this work, we focus on the core ranking problem within this semantic search framework. Specifically, we adopt a pointwise ranking structure that takes as input the textual information of the query and item (and additional user profile information, if appropriate), and predicts either the relevance between the query and item or the likelihood of user engagement. A product policy specifies how to grade the relevance of each (query, job) pair. The ranking system is then trained to optimize Normalized Discounted Cumulative Gain at the tenth position (NDCG@10) using policy-provided labels.
More concretely, given a query $q$ and a candidate item $j$, we aim to predict how relevant the pair $(q,j)$ are. In a basic cross-encoder ranking framework, one can be create a textual prompt, taking the form:
(1) $\text{prompt}(q,j)=\text{system prefix},q,j$
where the system prefix include chat template tags and instructions to the LLM to determine if the job posting $j$ is a good match to the query $q$. Given that the relevance problem studied here is a binary classification problem***In this work, we do not study setwise or listwise ranking frameworks., this textual prompt can be passed through an LLM with an appropriate classification output head, to estimate the relevance score. Particularly, the LLM (with a binary classification head) can estimate the probability distribution $(p_{\text{yes}},p_{\text{no}})$ where a higher value of $p_{\text{yes}}$ indicates a higher likelihood of relevance match:
(2) $p_{\text{yes}}(q,j)=\mathbb{P}(j~\text{is relevant to}~q)\in[0,1].$
LinkedIn Semantic Job Search
In this work, we focus on LinkedIn’s Semantic Job Search, which redefines job discovery through an AI-driven focus on intent rather than keywords. By interpreting natural language inputs, it captures users’ underlying goals and preferences, minimizing dependence on exact word matches between queries and job descriptions. This semantic understanding allows for more accurate and flexible retrieval, overcoming vocabulary gaps and delivering search results that align more closely with how members naturally express their career interests.
In practice, the prompt in (1) can be lengthy. In fact, in our semantic job search system, most of the prompt will be dominated with item description $j$, which has a median length of 900 tokens, and can take up to 2100 tokens. The lengthy prompts lead to unacceptable latency and throughput overhead, which makes deploying such a text-based system to production costly. To handle job search traffic spanning various product surfaces, our ranking system must process and score about 3.15 million items per second, making ranking efficiency absolutely critical.
Even though through previous summarization-solution(behdin2025scaling) we were able to reduce the item description length and rollout LLM ranking in limited-scale traffic, the prompt context length remains the major computational bottleneck and hard blocker for LinkedIn’s Semantic Job Search’s full deployment.
@@ 3. Architecture of MixLM
As the computational complexity of LLM inference scales quadratically with input length, reducing the number of tokens can substantially lower inference costs. In this work, we propose a text–embedding mix-interaction architecture that compresses item text description into an embedding vector using an encoder LLM. The resulting embeddings are then combined with other textual inputs and fed into a ranker LLM for output prediction. Because the item embeddings are precomputed and stored in a near-line cache, the number of tokens that must be processed during online inference is greatly reduced, leading to significant efficiency gains.
Detailed Structure
@@
$X_{E}$ $=\text{tokenize}([\text{prefix},j])\in\mathbb{V}^{T_{E}}$
where $T_{R}$ ($T_{E}$) is the length of $X_{R}$ ($X_{E}$). Note that $X_{E}$ does not depend on the query $q$ but only the item description $j$, and therefore, we can encode and store $X_{E}$ in an offline cache. We then let $h_{E}=F_{E}(X_{E})\in\mathbb{R}^{T_{E}\times H}$ to be the encoded item. To compress this item representation, we sample $h_{E}$ to a shorter sequence of length $T_{S}$ (specifically in MixLM, we use last N sampling), yielding $h_{S}=\mathrm{Samp}(F_{E}(X_{e}))\in\mathbb{R}^{T_{S}\times H}$. In general, one can consider several sampling techniques. In our work, for a give $T_{S}\geq 1$, we sample the $T_{S}$ embeddings in $h_{E}$ corresponding to the last $T_{S}$ input tokens. In particular, if $T_{S}=1$, we only take the embedding representation of the last input token.
Next, the input token embeddings of the prefix $X_{R}$ are computed via the ranker embedding layer, i.e., $h_{R}=g_{R}(X_{R})$. The two representations $h_{R},h_{S}$ are concatenated to form the combined embedding sequence $h_{P};h_{S}]\in\mathbb{R}^{(T_{R}+T_{S})\times H}$. The final “yes” probability is predicted by the ranker transformer:
(6) $p_{\text{yes}}(q,j;\Theta)=f_{R}\left([h_{P};h_{S}]\right)\in[0,1],$
@@
Since $h_{S}=\mathrm{Samp}(F_{E}(X_{e}))$ can be precomputed and stored in a near-line cache, the effective sequence length during online inference is reduced from $T_{R}+T_{E}$ to $T_{R}+T_{S}$. In practice, $T_{S}\ll T_{E}$, where we might take $T_{S}=1$ and $T_{E}$ can be up to 2100, representing orders of magnitudes of reduction in the prompt length. Furthermore, when ranking multiple items, $X_{R}$ is shared across all prompts that correspond to the same user and query. As the result, the Key–Value (KV) cache of $X_{R}$ of be used leading to further reduction of computational complexity.
@@ 4. Training
In this section, we describe the models, datasets and training recipes for the MixLM framework deployed to LinkedIn’s semantic job search system.
@@
4.1. Training Overview
Our MixLM framework requires us to train both the ranker and encoder LLMs. In particular, we require the encoded item embeddings available from the encoder model to align with the ranker model’s input embeddings, so the ranker model can make use of the encoded information. In general, as long as the ranker LLM’s hidden size is the same as encoder’s final output size (which allows us to concatenate embeddings in (6)), they can differ in architecture. However, to streamline the training process, we use models that have the same architecture for ranker and encoder, each with 0.6B parameters. We follow a three-stage training recipe for MixLM, summarized in Table 1. In stage I, we perform supervised fine-tuning on a 0.6B pretrained model using a domain-specific dataset (details to follow). The model resulting from this stage acts as the base model for our ranker model training. In stage II, using a purely text prompt, we train a ranker model using the labeled ground truth, which serves as a teacher model for our third training stage. This is motivated by our experience that presenting the model with complete textual item descriptions results in high ranking quality. In stage III, we co-train the ranker and the encoder models using custom loss functions that allow us to distill the text-based model from the second stage into a mixed input model. We use the model from the first training stage as the initial model checkpoint for the second stage and the ranker in the third stage, and a 0.6B General Text Embedding (GTE) model trained using contrastive training as the initial encoder checkpoint in the third stage.
@@ 4.2.1. Label Generation
We sampled query–item pairs from real user logs, and train a 7B LLM as our relevance judge to generate 5-point graded scores along several dimensions, together with rationales, such that it aligns with the product policy. The integer labels were subsequently normalized to continuous values in the range $[0,1]$ to create the ground truth labels $p^{*}_{\text{yes}}(q,j)$ for query-job pairs.
@@
We gather model’s chain-of-thought traces (wei2022chain), as well as the final matching score which we use in different stages of training.
@@ 4.2.2. Teacher Training Prompts in Stage II
               the job matches the query; no extra text.
@@ 4.2.3. Mixed Input Ranking Training Prompts in Stage III
               the job matches the query; no extra text.
@@
[CANDIDATE]  : <item embeddings>
yes or no ?
@@ 4.3.1. Stage I: Domain Reasoning Fine-Tuning
Concretely, we distill (hinton2014distilling) our 7B relevance judge into a 0.6B pretrained model, using KL divergence between the logits of the two models. As for the data, we randomly select 180180K samples following the prompt given in Section 4.2.1, together with the chain-of-thoughts responses of the relevance judge model.
@@ 4.3.2. Stage II: Ranking Teacher Training
Specifically, starting from the checkpoint available from our first training stage, we perform Supervised Fine-Tuning (SFT) using the input prompt discussed in Section 4.2.2. We generate 10.910.9M examples based on the prompt provided above, and normalize the integer matching scores, available from the 7B judge, into ground truth probabilities $(p^{*}_{\text{yes}},p^{*}_{\text{no}})$. The SFT is then performed using a KL divergence loss as in (LABEL:eq:kl_loss). We use $(\hat{p}_{\text{yes}},\hat{p}_{\text{no}})$ to denote the output probability distribution from the final checkpoint for this training stage.
@@ 4.3.3. Stage III: Joint Encoder-Ranking Training
This loss function is inspired by knowledge distillation (hinton2014distilling), where knowledge from a (typically larger) teacher model is transferred to a (usually smaller) student model. In contrast, our student model, MixLM, which consists of an encoder and a ranker, can in fact be larger than the teacher model. As discussed in Section 6, the teacher operating on pure text is easier to train to provide high-quality predictions that are cleaner than the original dataset labels. When training MixLM, the teacher label helps reducing gradient variance and facilitates better convergence.
@@
Like multi-modal language models, such as LLaVA (liu2023visual) and BLIP-2 (li2023blip), the goal of Stage III is to align the input space of the ranker LLM and the embedding space of the encoder LLM. This is challenging because 1) unlike the discrete vocabulary of text, item embeddings have an infinitely large continuous space, resulting in higher sample complexity; 2) the embeddings encode much more complex and entangled information than text tokens, requiring greater training effort to preserve and decode such information; and 3) item embeddings have never been encountered during foundational pretraining.
@@
This approach mirrors the multi-stage training strategies successfully employed in multi-modal LMs, where phased training with different optimization objectives in each phase improves both alignment quality and downstream task performance. During the training, we employ a standard AdamW optimizer, combined with a cosine annealing scheduler with linear warmup.
4.4. Training Efficiency
During the ranking training phase, we co-train the encoder LLM and the ranker LLM end-to-end. We deploy PyTorch’s Fully Sharded Data Parallel (zhao2023fsdp) to shard the decoder parameters, gradients, and optimizer states across 8 nodes (each with 8 NVIDIA H100 GPUs). Within each node, GPUs communicate via NVLink, and inter-node communication is via InfiniBand.
To reduce communication overhead, we enable layer-wise prefetching: before computing the next forward layer, we prefetch its parameters; during the backward pass, we prefetch gradients of upcoming layers. To further reduce memory footprint, we incorporate Liger Kernel (hsu2024liger), a set of optimized Triton kernels (e.g. fused RMSNorm, RoPE, SwiGLU, fused CrossEntropy) that deliver higher throughput and lower memory usage.
In our setup, the total compute cost to train on 70B tokens under the ranking objective is approximately 700 H100 GPU-hours.
5. System Implementation
5.1. Online System
The online inference architecture adopts a two-layer design consisting of a feature generation layer and a serving layer (see Figure 3 for an overview). The feature generation layer functions as a hybrid system that combines offline refresh with near-line ingestion and generation to maintain comprehensive and up-to-date feature coverage. The offline pipeline, orchestrated through Flyte, periodically performs large-scale GPU-based inference to regenerate the full corpus of text and embedding features. In parallel, the near-line component continuously ingests real-time updates reflecting changes in the population of scoring candidates and generates corresponding features through GPU inference. Because near-line processing is not latency-sensitive, GPU resources can be utilized with high throughput and efficiency. All generated features are persistent in a high-throughput distributed storage system, ensuring consistency, low-latency availability for downstream serving.
The serving layer performs real-time inference by retrieving precomputed features and assembling mixed text–embedding prompts that integrate user queries with candidate scoring items. Prior to dispatching requests to the SGLang inference engine (sglang) for GPU-based scoring, several system-level optimizations—including distributed caching, latency reduction mechanisms, and scalable serving infrastructure—are applied to minimize feature retrieval overhead and improve serving efficiency. This layered architecture enables effective reuse of precomputed representations while preserving responsiveness across both near-line and real-time applications. Empirical evaluations demonstrate that these design choices significantly enhance inference efficiency, GPU utilization, and system scalability, supporting production-scale deployment of large language model–based ranking systems.
Figure 3. Online Serving Architecture for LinkedIn’s Semantic Job Search
5.2. Inference Engine
5.2.1. Text-Embedding Mixed Input

### v2

@@ Abstract.
[…]
Semantic Search, Large Language Models, Learning to Rank
copyright: acmlicensedjournalyear: 2025conference: Preprint; 2025; Preprintccs: Computing methodologies Information retrievalccs: Computing methodologies Natural language processing11footnotetext: Equal key contributors.22footnotetext: Work done while at LinkedIn.33footnotetext: Correspondence authors.
@@ 1. Introduction
Large Language Models (LLMs) have been widely adopted in industrial applications due to their strong generative and predictive capabilities. In search and recommender systems, LLMs deliver substantial relevance improvements by capturing subtle linguistic cues and conceptual relationships that traditional lexical methods often miss (Qin et al., 2024; Meng et al., 2025; Hou et al., 2024; Wu et al., 2024a). When used as relevance judges or rankers, LLMs enable semantic retrieval and ranking that better aligns with user intent.
Despite these strengths, deploying LLMs in industrial search and recommendation systems remains challenging due to their high computational cost, particularly when processing long text inputs under strict latency and throughput constraints. A common cross-encoder approach forms prompts that combine user queries with candidate item descriptions and asks the LLM to assess relevance (Qin et al., 2024; Meng et al., 2025; Ma et al., 2024; Pradeep et al., 2023). Such prompts often contain thousands of tokens, resulting in heavy prefill-dominated inference and quadratic attention costs with respect to context length. In practice, this forces systems to either replace LLMs with smaller models or drop informative features from online serving pipelines (Dey et al., 2025; Wang et al., 2024; Shang et al., 2025). Reducing input context length while preserving semantic expressiveness is therefore a central challenge for LLM-based ranking.
@@
In this work, we introduce MixLM, an LLM-based ranking framework designed to substantially reduce input context length while retaining the semantic richness of full-text LLM rankers. MixLM compresses each candidate document—often thousands of tokens long—into a small set of learned embedding tokens using an encoder LLM. These embeddings are stored in a nearline cache and, at inference time, are mixed with the user’s natural-language query and auxiliary text and processed by a ranker LLM. By replacing long document text with a handful of embedding tokens, MixLM reduces computational complexity by orders of magnitude while preserving rich query–document interactions.
@@
Beyond the modeling formulation, we describe the end-to-end training and deployment of MixLM in LinkedIn’s production semantic job search system. We introduce a joint training pipeline that optimizes both encoder and ranker LLMs and propose distillation losses that transfer knowledge from a strong full-text LLM ranker into the mixed-interaction architecture. We also analyze scaling behavior with respect to training data size and the number of embedding tokens per item.
@@
From a systems perspective, we present a detailed analysis of MixLM’s impact on serving infrastructure. Context compression via embedding tokens enables substantially higher ranking throughput, which is further amplified by nearline caching and amortized prefill computation. Together, these optimizations compound to deliver large efficiency gains for mixed-input LLM serving.
@@
We validate MixLM at industrial scale through deployment in LinkedIn’s AI-driven Job Search. Under the same latency budget, MixLM:
[…]
@@
These efficiency gains enabled full-traffic deployment of LLM-based ranking and resulted in a 0.47% lift in Daily Active Users (DAU) in a large-scale online A/B experiment.
Summary of Contributions.
  • We introduce MixLM, a mixed-input LLM ranking framework that replaces long item text with compact embedding tokens to reduce prompt length.
  • We propose an end-to-end training pipeline with distillation losses that align mixed-input ranking with full-text LLM behavior.
  • We present a production-ready serving stack for mixed-input LLM ranking, achieving order-of-magnitude improvements in GPU throughput.
  • We share practical lessons from deploying MixLM in LinkedIn’s semantic job search at full production scale.
Together, these contributions demonstrate a practical and scalable path for deploying LLM-based ranking systems that combine deep semantic understanding with real-time efficiency.
@@ 2. Semantic Search
Semantic search retrieves and ranks items based on semantic similarity rather than lexical overlap. A standard pipeline consists of dense representation learning, large-scale vector retrieval (e.g., GPU-accelerated exhaustive search), high-capacity reranking to model query–item interactions, and downstream business logic such as personalization and filtering.
We focus on the ranking component and adopt a pointwise formulation in which a model consumes the query, candidate item, and optional user context to predict relevance or user engagement. Relevance labels are defined by product policy, and models are trained to optimize NDCG@10 using policy supervision.
Given a query $q$ and candidate item $i$, a cross-encoder ranker constructs a single textual prompt
(1) $\text{prompt}(q,i)=\text{system prefix},q,i,$
where the system prefix includes instructions guiding the LLM to assess semantic match.
Since relevance is treated as a binary classification task***We do not consider setwise or listwise ranking in this work., the prompt is processed by an LLM with a classification head that estimates
(2) $p_{\text{yes}}(q,i)=\mathbb{P}(i~\text{is relevant to}~q)\in[0,1],$
with higher $p_{\text{yes}}$ indicating stronger relevance.
@@ 3. Architecture of MixLM
As the computational complexity of LLM inference scales with input length, reducing the number of tokens can substantially lower inference costs. In this work, we propose a text–embedding mix-interaction architecture that compresses item text description into an embedding vector using an encoder LLM. The resulting embeddings are then combined with other textual inputs and fed into a ranker LLM for output prediction. Because the item embeddings are precomputed and stored in a near-line cache, the number of tokens that must be processed during online inference is greatly reduced, leading to significant efficiency gains.
Detailed Architecture
@@
$X_{E}$ $=\text{tokenize}([i])\in\mathbb{V}^{T_{E}}$
where $T_{R},T_{E}$ are the lengths of $X_{R},X_{E}$. Note that $X_{E}$ does not depend on the query $q$ but only the item description $j$, and therefore, we can encode and store $X_{E}$ nearline/offline. We then let $h_{E}=f_{E}(X_{E})\in\mathbb{R}^{T_{E}\times H}$ to be the encoded item. To compress this item representation, we sample $h_{E}$ to a shorter sequence of length $T_{S}$ (specifically in MixLM, we use last N sampling), yielding $h_{S}=\mathrm{Samp}(f_{E}(X_{E}))\in\mathbb{R}^{T_{S}\times H}$. In general, one can consider several sampling techniques. In our work, for a give $T_{S}\geq 1$, we sample the $T_{S}$ embeddings in $h_{E}$ corresponding to the last $T_{S}$ input tokens. In particular, if $T_{S}=1$, we only take the embedding representation of the last input token.
Next, the input token embeddings of the prefix $X_{R}$ are computed via the ranker embedding layer, i.e., $h_{R}=g_{R}(X_{R})$. The two representations $h_{R},h_{S}$ are concatenated to form the combined embedding sequence $h_{R};h_{S};h_{EOS}]\in\mathbb{R}^{(T_{R}+T_{S}+1)\times H}$, followed by a special end of sentence token. The final “yes” probability is predicted by the ranker transformer:
(6) $p_{\text{yes}}(q,i;\Theta)=f_{R}\left([h_{R};h_{S};h_{EOS}]\right)\in[0,1],$
@@
Since $h_{S}=\mathrm{Samp}(F_{E}(X_{e}))$ can be precomputed and stored in a near-line cache, the effective sequence length during online inference is reduced from to $T_{R}+T_{S}+1$. In practice, $T_{S}\ll T_{E}$, where we might take $T_{S}=1$ and $T_{E}$ is 2100 in p99, representing orders of magnitudes of reduction in the prompt length. Furthermore, when ranking multiple items, $X_{R}$ is shared across all prompts that correspond to the same user and query. As the result, the Key–Value (KV) cache of $X_{R}$ of be used leading to further reduction of computational complexity.
@@ 4. Training
In this section, we describe the models, datasets, and training pipeline used to deploy the MixLM framework in LinkedIn’s semantic job search system. MixLM requires jointly training a ranker LLM and an encoder LLM, with the key requirement that the encoder’s output embeddings align with the ranker’s input embedding space so that encoded item representations can be consumed directly by the ranker (via concatenation in Eq. (6)). As long as the ranker’s hidden size matches the encoder’s output dimension, the two models may differ architecturally. For simplicity and training efficiency, however, we use the same 0.6B-parameter architecture for both the ranker and encoder.
We adopt a three-stage training recipe for MixLM, summarized in Table 1. In Stage I, we perform supervised fine-tuning of a 0.6B pretrained model on a domain-specific dataset; the resulting checkpoint serves as the base initialization for the ranker. In Stage II, we train a pure-text ranker using full item descriptions and labeled ground truth. This model achieves high ranking quality and is treated as a teacher for subsequent distillation. In Stage III, we co-train the ranker and encoder using custom loss functions that distill knowledge from the text-based teacher into the mixed-input architecture. The ranker is initialized from the Stage I checkpoint, while the encoder is initialized from a 0.6B General Text Embedding (GTE) model trained with contrastive learning.
@@ 4.1.1. Label Generation
We sampled query–item pairs from real user logs, and train a 7B LLM as our relevance judge to generate 5-point graded scores along several dimensions, together with rationales, such that it aligns with the product policy. The integer labels were subsequently normalized to continuous values in the range $[0,1]$ to create the ground truth labels $p^{*}_{\text{yes}}(q,j)$ for query-item pairs.
@@
We gather model’s chain-of-thought traces (Wei et al., 2022), as well as the final matching score which we use in different stages of training.
@@ 4.1.2. Teacher Training Prompts in Stage II
               the item matches the query; no extra text.
@@ 4.1.3. Mixed Input Ranking Training Prompts in Stage III
               the item matches the query; no extra text.
@@
[CANDIDATE]  : <item embeddings> <EOS>
@@ 4.2.1. Stage I: Domain Reasoning Fine-Tuning
Concretely, we distill (Hinton, 2014) our 7B relevance judge into a 0.6B pretrained model, using KL divergence between the logits of the two models. As for the data, we randomly select 180180K samples following the prompt given in Section 4.1.1, together with the chain-of-thoughts responses of the relevance judge model.
@@ 4.2.2. Stage II: Ranking Teacher Training
Specifically, starting from the checkpoint available from our first training stage, we perform Supervised Fine-Tuning (SFT) using the input prompt discussed in Section 4.1.2. We generate 10.910.9M examples based on the prompt provided above, and normalize the integer matching scores, available from the 7B judge, into ground truth probabilities $(p^{*}_{\text{yes}},p^{*}_{\text{no}})$. The SFT is then performed using a KL divergence loss. We use $(\hat{p}_{\text{yes}},\hat{p}_{\text{no}})$ to denote the output probability distribution from the final checkpoint for this training stage.
@@ 4.2.3. Stage III: Joint Encoder-Ranking Training
This loss function is inspired by knowledge distillation (Hinton, 2014), where knowledge from a (typically larger) teacher model is transferred to a (usually smaller) student model. In contrast, our student model, MixLM, which consists of an encoder and a ranker, can in fact be larger than the teacher model. As discussed in Section 6, the teacher operating on pure text is easier to train to provide high-quality predictions that are cleaner than the original dataset labels. When training MixLM, the teacher label helps reducing gradient variance and facilitates better convergence.
@@
5. Inference Engine Optimization
5.1. Text–Embedding Mixed Input